OpenStudy (anonymous):

A car dealership interviewed every driver who came onto their lot on Tuesday. This produced the following results: there were 68 people who owned trucks and 121 people who owned cars. If 35 people owned both a truck and a car and 15 people owned neither, how many people were interviewed?

OpenStudy (shane_b):

Add up those who owned trucks and cars: \[121+68=189\]Subtract off the number of people who owned both:\[189-35=154\]Finally, add in the number of people who owned neither:\[154+15=169\]